ADNAN MUHAMMED

Grade 12,

A particle moving with a uniform acceleration travels 24m and 64m in the first two consecutive intervals of 4 seconds each. Its initial velocity is?

A particle moving with a uniform acceleration travels 24m and 64m in the first two consecutive intervals of 4 seconds each. Its initial velocity is?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Fatima
 
s = ut + 1/2 at2
so distance travelled in first 4 seconds = 
24 = u(4) + 1/2 a(42)  =>24 = 4u + 8a - (i)
​now total distance travelled in 8s = 
(24+64) = u(8) + 1/2 a(82) =>88 = 8u + 32a -(ii)
multiplying equation (i) by 4 we have,
96 = 16u + 32a-(iii)
subtracting equation (ii) from equation (iii) we have,
8 = 8u
​=> u = 1m/s
 
Regards
Arun